OpenStudy (anonymous):

The path of a ball is given by y = -1/20x^2 + 3x + 5, where y is the height (in feet) of the ball and x is the horizontal distance (in feet) from where the ball was thrown. Find the maximum height of the ball. Which number determines the height at which the ball was thrown? Does changing this value change the coordinates of the maximum height of the ball? Explain.

OpenStudy (mertsj):

y'=-.1x+3 =0 x=30 y=-1/20(30)^2+3(30)+5 y=50= maximum height of the ball.

OpenStudy (anonymous):

Thank you, how did you get this answer?

OpenStudy (mertsj):

y'=-.1x+3 =0 x=30 y=-1/20(30)^2+3(30)+5 y=50= maximum height of the ball.

OpenStudy (anonymous):

also do you know the answer to the second part - Does changing this value change the coordinates of the maximum height of the ball? Explain.

OpenStudy (mertsj):

I took the first derivative and set it equal to 0 to find the x coordinate of the maximum point. Having found the x value of the maximum point, I plugged it into the original equation to find the y coordinate which is the maximum

OpenStudy (anonymous):

wow, u r amazing, thank you so very much, i wish i understood this stuff

OpenStudy (anonymous):

do u know the 2nd part of the problem?

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Can anyone help me please?

OpenStudy (mertsj):

The constant term determines the initial height of the ball. Yes. If you change that number it does change the maximum height of the ball because that number is used to find the y coordinate of the vertex which is the maximum height of the ball.

OpenStudy (anonymous):

ok, thank you...you are awesome.

OpenStudy (mertsj):

yw